OpenStudy (anonymous):

solve the equation and check solution z^2+6z=-5

sam (.sam.):

z^(2)+6z=-5 To set the left-hand side of the equation equal to 0, move all the expressions to the left-hand side. z^(2)+6z+5=0 In this problem 5*1=5 and 5+1=6, so insert 5 as the right hand term of one factor and 1 as the right-hand term of the other factor. (z+5)(z+1)=0 z=-5,-1